# Active nematic pumps

### v1

@@ Abstract
Microfluidics involves the manipulation of flows at the microscale, typically requiring external power sources to generate pressure gradients. Alternatively, harnessing flows from active fluids, which are usually chaotic, has been proposed as a paradigm for the development of micro-machines. Here, by combining experimental realizations and simulations, we demonstrate that the addition of triangular-shaped inclusions into an active nematic gel can locally break the fore-aft symmetry of active turbulence and stabilize flow fields with self-pumping capabilities. The proposed strategy has enabled us to generate wall-free and self-powered microfluidic systems capable of both cargo transport and mixing along with the downstream flow. We analyze the performance of these active pumps, both isolated and within cooperative ensembles in terms of their output velocity and hydrostatic pressure buildup. Finally, we demonstrate strategies to incorporate them into specifically designed microfluidic platforms to advantageously tailor the geometry of active flows. Our results reveal new possibilities for leveraging the self-organized mechanodynamics of active fluids.
@@ I Introduction
Figure 1: Active pumping with a square array of triangular inclusions. a An array of hydrogel columns with triangular cross section is photopolymerized in-situ in the aqueous phase of an AN flow cell. Scale bar 100 $\mu$m. b Fluorescence image showing the adaptation of the AN to the presence of the obstacle array (see Movie Active nematic pumps). c Simulated system (see Movie Active nematic pumps). The instantaneous director field is rendered using the Line Integral Convolution filter. Diagrams d and f correspond to the analysis of the experimental flow field, while e and g correspond to the analysis of the simulated flow field. Experimental images correspond to the central part of a lattice containing 6 rows and 9 columns in total, while the simulated system, also only partially shown, contains 6 rows and 30 columns with periodic boundary conditions in both directions. Time-averaged velocity (d, e) and vorticity (f, g) fields. $v_{r}$ is the average speed and, in the simulations, $\omega_{r}=v_{r}/D_{def}$, where $D_{def}$ is the average distance between defects, all measured in the absence of obstacles. h Simulated average horizontal flow velocity relative to $v_{r}$ as a function of the size of the equilateral triangle, $H$, for three different normalized activity coefficients, $\zeta D_{def}/(v_{r}\eta)$, and for a lattice periodicity $L=128$ l.u. (or $\rm 2.3\,D_{def}$). Here, $\eta$ is the two-dimensional active fluid viscosity. Simulations are performed in one lattice cell, using periodic boundary conditions. Data are collapsed by rescaling $H$ with $D_{def}$. i Diagram of the normalized average horizontal velocity, $v_{x}/v_{r}$, as a function of the normalized triangle size, $H/D_{def}$, and triangle size relative to cell size, $H/L$. Simulations are performed with $\zeta D_{def}/(v_{r}\eta)=7.84$.
@@
Here, we address this gap by demonstrating that inclusions with well-defined geometrical asymmetry can lead to the phenomenon of active pumping. In particular, we take advantage of a recent protocol to polymerize biocompatible hydrogel structures directly into a fully-developed active nematic gel [29] to imprint lattices of triangular inclusions that, for sizes comparable to the intrinsic active length scale, lead to hydrostatic pressure buildup and enhanced unidirectional flows. Combining experiments and numerical simulations we show that the observed phenomenon is a cooperative effect emerging from a lattice of counter-rotating vortices that result from the presence of obstacles with broken fore-aft symmetry within the active nematic fluid. By characterizing the interplay between flow patterns and pressure gradients generated within the active nematic fluid, we demonstrate that arrays of triangular inclusions can lead to efficient active nematic pumping through the creation of virtual flow channels. The absence of physical confinement walls makes this system an ideal candidate for large-scale active microfluidic transport and mixing of colloidal cargo, or to locally adjust the geometry of the active fluid [30].
@@
Figure 2: Active pumps in series and in parallel. a Simulated study of the performance of pumps in series. The number of pumps acting in parallel (number of flow channels) is kept constant at 3, while the number of columns, $N$ is varied. The yellow square represents a single pump in the region between four triangles. The size of the simulated domain, $36\times 6$ cells (or $\rm 82.3\times 13.7\,D_{def}$), and the simulation conditions, $H=1.25D_{def}$, $L=2.3D_{def}$, $\zeta D_{def}/(v_{r}\eta)=7.84$ and closed box with no-slip boundary conditions, are the same in all cases. b Velocity field along the central lines between rows of triangles (averaged for the three rows and with time) for different numbers of pumps in series, including $N=0$ (closed box without triangles) for comparison. The orange dashed vertical line indicates the exit of the pump region. c For the same simulations as in (b), we monitor the pressure field. Here we plot the normalized pressure profile, averaged for the three rows and with time, and normalized with $p_{r}=\rho_{0}v_{r}^{2}$, where $\rho_{0}$ is the average density. d Average normalized horizontal velocity at the output of the pump system for different numbers of pumps in series. The inset indicates the corresponding pressure difference across the pump system. The error bars represent the standard deviation for different random initial conditions. e System of pumps in parallel. The number of pumps acting in series is kept constant at 9 (10 columns of triangles), while the number of rows, $N$, is varied. f Average flow rate at the output for different numbers of pumps in parallel, relative to the reference flow rate $Q_{r}=v_{r}D_{def}$. The inset depicts the pressure difference across the pump system.
@@ II.1 Creation of active nematic pumps
We photopolymerize square lattices of hydrogel columns with triangular cross section inside an AN flow cell [29]. The columns are attached to the hydrophilic glass plate and extend across the AN towards the thin oil layer on the complementary hydrophobic plate (see Fig. 1a and Methods). As a result, a lattice of rigid triangular obstacles is suddenly embedded within the AN, thus forcing the adaptation of its orientational and velocity fields to the new in-plane boundary conditions (Fig. 1b).
@@
Both experiments and continuum hydrodynamic simulations (Fig. 1c; see Methods) show that the active flow patterns organize around the array of obstacles leading to the formation of virtual flow channels. The velocity field is oriented, on average, along the space between rows of obstacles, antiparallel to the orientation of the triangles in the array (left-to-right in the images), as shown in Fig. 1d and Fig. 1e. This is different from active nematic flows confined in channels, which can exhibit net transport of the normally turbulent-like flows, but without a predefined flow polarity [17, 22]. In particular, we observe that active flows organize as a pair of counter rotating vortices on both sides of the frontal tip of each triangle. This is clearly highlighted in Fig. 1f and g, where we represent the average vorticity field of the active flows. Our observations are consistent with earlier simulations that reported a steady vortex lattice forming in AN confined in an array of astroid-shaped obstacles [31]. In our case, the flows emerging from counter rotating vortex pairs leads to net transport in a direction that is predefined by the orientation of the obstacles in the array. Inside the virtual channels, flow speed is non monotonous, being higher between vortex pairs. Overall, this leads to a unidirectional pumping mechanism. (Fig. Active nematic pumps). A similar flow rectification effect has been recently proposed using numerical simulations with half-astroid-shaped obstacles [32], consistent with the expected ratchet effect in the presence of asymmetric obstacles.
@@
The reported pumping mechanism depends on the number, size, shape, and spatial distribution of the triangles in the array. A stable vortex lattice emerges only if at least 4 inclusions are present (Fig. Active nematic pumps). We consider a 2 by 2 square lattice of triangular inclusions to be the minimal system to achieve active pumping. In Fig. 1, the size and spacing of the triangles are chosen to be similar to the intrinsic length scale of the AN, $D_{def}$, which we have determined to be of the order of 70 $\mu$m from the average spacing between defects in the absence of obstacles (see early stages of Movie S1). Using numerical simulations, we have studied the effect of these two parameters.
@@
In Fig. 1h, we show the average normalized horizontal velocity as a function of the normalized size of the triangles, $H/D_{def}$, for a fixed spacing $L=2.3D_{def}$. Note that, in the absence of AN pumping effects, the flows are turbulent-like and therefore, the net velocity is zero in any direction. We observe a quick increase of the pumping efficiency with obstacle size, until $H\simeq D_{def}$, slowly decreasing for larger obstacles. We have performed simulations for different activity coefficients (related to the concentration of adenosine triphosphate in experiments), and find that the results remain invariant as the velocity and spatial dimensions are rescaled with the intrinsic velocity and active length scale in the absence of obstacles at each activity [33, 20].
@@
We next use simulations to explore, for a constant activity, the optimal geometry of the square lattice of triangles, in terms of their size and spacing. The resulting phase diagram, shown in Fig. 1i, indicates that the optimal arrangement corresponds to $H\simeq D_{def}$ and $L\simeq 2D_{def}$, which we have used in the experiments and simulations in Fig. 1, and will be used henceforth by default. Our experimental technique allows to fabricate more compact active pump arrangements, as shown in Fig. Active nematic pumps, thus demonstrating the capabilities for miniaturization of these devices.
@@ II.2 Active pump systems
After determining the optimal geometric features of basic AN pumps, we consider the design of more intrincate microfluidic systems. In particular, we have studied the performance of square lattices of triangles (Fig. 2) for different number of columns (active pumps in series, Fig. 2a-d) and rows (active pumps in parallel, Fig. 2e,f). We observed that, by concatenating AN pumping units, the average speed increases with the number of columns (Fig. 2b), until reaching a maximum value at about 10 pumps in series (Fig. 2d).
In parallel, we also consider the concurrent profiles of the hydrostatic pressure within the active fluid along the flow channels, and we observe that it increases steadily with the number of pumping elements in series, even when the average speed has reached its saturation value (Fig. 2c and inset in Fig. 2d). In our active pump system, pressure gradients are positively correlated with the flow velocity. This is contrary to flow through a microfluidic channel or through a porous medium, where velocity is negatively correlated with the driving pressure gradients, even in the case of pressurized active fluids [34]. In classical hydraulics, a system of ideal pumps in series yields a flow velocity that does not change with the number of pumps and a pressure that builds up linearly with the number of elements [35]. Interestingly, our active pumps behave as ideal hydraulic pumps when the number of elements in series is larger than 10, as observed in Fig. 2d.
@@
Figure 4: Microfluidic applications of active nematic pumping. a Experiments where the output of an active pump system is injected into a channel. Triangular obstacles and channel boundaries are overlaid on the map of the velocity field. Scale bar 100 $\mu$m. b Horizontal velocity (averaged over the central $50\,\mu$m along the channel) relative to $v_{r}$. c Simulation of an active pump system injecting into a channel of width $L_{c}=2.3D_{def}$ and activity $\zeta D_{def}/(v_{r}\eta)=7.84$. We perform simulations inside a closed box of dimensions $24\times 6$ cells (or $\rm 54.9\times 13.7\,D_{def}$), with no-slip boundary conditions on the walls. d Velocity field along the central part of the channel, averaged in time and for the different initial conditions in simulations, for three different normalized channel widths $L_{c}/D_{def}$. e Velocity field inside the channel with and without the pump system for different channel widths in simulations. The error bars are obtained from the different initial conditions. f Simulations of larger scale flows guided by a lattice of triangles. Clockwise flow circulation is guided by the active pump system. The white and the grey trajectories illustrate the path followed by two tracers and the white rectangle indicates the path used to calculate the circulation. The system, with dimensions $16\times 10$ cells (or $\rm 36.6\times 18.3\,D_{def}$), is in a closed box with solid walls at the borders. g Circulation, $C=\oint\mathbf{u}\cdot\mathbf{dl}$, vs. time along the rectangular circuit in the simulations in f. h Experimental time-averaged velocity field in the presence of a closed circuit of elementary active pumps. Scale bar $100\mu$m.
Next, we analyze the performance of active pumps in parallel, as shown in Fig. 2e. In this case, the setup includes a fixed number of 10 columns of triangles (9 pumps in series), and we study the performance of the system when changing the number of rows. The output velocity is the one expected from 9 pumps in series, so we will characterize each pump system with its average flow rate, which we find to be proportional to the number of elements in parallel, just as it would be for a system of ideal hydraulic pumps (Fig. 2f). We have also monitored the average pressure drop across the system, which we find to be constant, independent of the number of elements in parallel. These observations show that neighboring pumping elements along the same column operate independently, as a system of ideal pumps, despite the fact that the separation between adjacent flow channels is permeable. This favors not only net transport but also matter exchange, as we show in the next section.

### v2

@@ Abstract
Abstract. Microfluidics involves the manipulation of flows at the microscale, typically requiring external power sources to generate pressure gradients. Alternatively, harnessing flows from active fluids, which are usually chaotic, has been proposed as a paradigm for the development of micro-machines. Here, by combining experimental realizations and simulations, we demonstrate that the addition of triangular-shaped obstacles into an active nematic gel can locally break the fore-aft symmetry of active turbulence and stabilize flow fields with self-pumping capabilities. The proposed strategy has enabled us to generate wall-free and self-powered microfluidic systems capable of both cargo transport and mixing along with the downstream flow. We analyze the performance of these active pumps, both isolated and within cooperative ensembles in terms of their output velocity and hydrostatic pressure buildup. Finally, we demonstrate strategies to incorporate them into specifically designed microfluidic platforms to advantageously tailor the geometry of active flows. Our results reveal new possibilities for leveraging the self-organized mechanodynamics of active fluids.
Significance statement. Lab-on-a-chip devices allow multi-sample chemical analysis and complex in-situ microfabrication in an integrated platform where fluids are propelled and controlled at submillimeter scales. While these systems are typically hampered by the need to rely on external power sources, we demonstrate here an alternative strategy using active fluids, composed of self-propelled units that normally develop disordered flows. We work with an active nematic fluid where we embed lattices of triangular hydrogel micro-columns, which transform the turbulence into controllable flow patterns, effectively harnessing the local activity and leading to the emergence of active pumps that can be arbitrarily combined to internally adjust the pressure or flow rate in an integrated microfluidic device.
@@ I Introduction
Figure 1: Active pumping with a square array of triangular obstacles. a An array of hydrogel columns with triangular cross section is photopolymerized in-situ in the aqueous phase of an AN flow cell. Scale bar 100 $\mu$m. b Fluorescence image showing the adaptation of the AN to the presence of the obstacle array (see Movie Active nematic pumps). c Simulated system (see Movie Active nematic pumps). The instantaneous director field is rendered using the Line Integral Convolution filter. Diagrams d and f correspond to the analysis of the experimental flow field, while e and g correspond to the analysis of the simulated flow field. Experimental images correspond to the central part of a lattice containing 6 rows and 9 columns in total, while the simulated system, also only partially shown, contains 6 rows and 30 columns with periodic boundary conditions in both directions. Time-averaged velocity (d, e) and vorticity (f, g) fields. $v_{r}$ is the average speed and, in the simulations, $\omega_{r}=v_{r}/D_{def}$, where $D_{def}$ is the average distance between defects, all measured in the absence of obstacles. h-j Average active nematic fields around triangular obstacles. Left column: experiments; right column: simulations. Conditions are the same as in the other panels in this figure. The field of view in simulations is exactly one cell in the periodic lattice, while in experiments the field of view is slightly larger than the unit cell. h Scalar nematic order parameter, $S$, and director field. i Topological charge density, $\sigma$, and director field. j Vorticity field with hotspots from the topological charge density map highlighted.
Figure 2: Optimal active pump geometry. a Simulated average horizontal flow velocity relative to $v_{r}$ as a function of the size of the equilateral triangle, $H$, for three different normalized activity coefficients, $\zeta D_{def}/(v_{r}\eta)$, and for $H/L=0.55$. Here, $\eta$ is the two-dimensional active fluid viscosity. Simulations are performed in one lattice cell, using periodic boundary conditions. Data are collapsed by rescaling $H$ with $D_{def}$. b Diagram of the normalized average horizontal velocity, $v_{x}/v_{r}$, constructed from simulations, as a function of the normalized triangle size, $H/D_{def}$, and triangle size relative to cell size, $H/L$. Simulations are performed with $\zeta D_{def}/(v_{r}\eta)=7.84$.
@@
Here, we address this gap by demonstrating that obstacles with well-defined geometrical asymmetry can lead to the phenomenon of active pumping. In particular, we take advantage of a recent protocol to polymerize biocompatible hydrogel structures directly into a fully-developed active nematic gel [29] to imprint lattices of triangular obstacles that, for sizes comparable to the intrinsic active length scale, lead to hydrostatic pressure buildup and enhanced unidirectional flows. Combining experiments and numerical simulations we show that the observed phenomenon is a cooperative effect emerging from a lattice of counter-rotating vortices that result from the presence of obstacles with broken fore-aft symmetry within the active nematic fluid. By characterizing the interplay between flow patterns and pressure gradients generated within the active nematic fluid, we demonstrate that arrays of triangular obstacles can lead to efficient active nematic pumping through the creation of virtual flow channels. The absence of physical confinement walls makes this system an ideal candidate for large-scale active microfluidic transport and mixing of colloidal cargo, or to locally adjust the geometry of the active fluid [30]. Confinement-free effects also distinguish our scenario from that of the pumping of a threshold-less active nematic suggested by Green et al. [31]. In this case, flows are triggered to balance the force density created by geometrical distortions.
@@
Figure 3: Active pumps systems. a Simulated study of the performance of an array of pumps as we change the number of columns, $M$. The number of rows is kept constant at 4 (3 flow channels). The yellow square represents a single pump in the region between four triangles. The size of the simulated domain is $83.2\,D_{def}\times 13.7\,D_{def}$. The simulation conditions are $H=1.25D_{def}$, $L=2.3D_{def}$, $\zeta D_{def}/(v_{r}\eta)=7.84$ and closed box with no-slip boundary conditions, are the same in all cases. b Velocity field along the central lines between rows of triangles (averaged for the three rows and with time) for different number of columns, including $M=0$ (closed box without triangles) for comparison. The orange dashed vertical line indicates the exit of the pump region. c For the same simulations as in (b), we monitor the pressure field for three different values of $M$. Here we plot the normalized pressure profile, averaged for the three rows and with time, and normalized with $p_{r}=\rho_{0}v_{r}^{2}$, where $\rho_{0}$ is the average density. d Average normalized horizontal velocity at the output of the pump array for different number of columns. The inset indicates the corresponding pressure difference across the pump system. The error bars represent the standard deviation for different random initial conditions. e Array of pumps where we keep the number of columns fixed at 10 and we vary the number of rows, $N$. The size of the simulated domain is $\rm 37.7\,D_{def}\times 43.4\,D_{def}$. The size of the unit cell and the simulation conditions are the same as above. f Average flow rate at the output for different number of rows, relative to the reference flow rate $Q_{r}=v_{r}D_{def}$. The inset depicts the pressure difference across the pump system.
@@ II.1 Creation of active nematic pumps
We photopolymerize square lattices of hydrogel columns with triangular cross section inside an AN flow cell [29]. The columns are attached to the hydrophilic glass plate and extend across the AN towards the thin oil layer on the complementary hydrophobic plate (see Fig. 1a and Methods). As a result, a lattice of rigid triangular obstacles is suddenly embedded within an AN layer in the active turbulent regime. Such intervention prompts the adaptation of its orientational and velocity fields to the new in-plane boundary conditions (Fig. 1b). Upon polymerization, filaments in irradiated regions are frozen in non-specified orientations. This seems to have no impact either on the surrounding orientational or velocity fields of the embedded AN, which organize similarly around all obstacles.
@@
Both experiments and continuum hydrodynamic simulations (Fig. 1c; see Methods) show that the active flow patterns organize around the array of obstacles leading to the formation of virtual flow channels. The velocity field is oriented, on average, along the space between rows of obstacles, antiparallel to the orientation of the triangles in the array (left-to-right in the images), as shown in Fig. 1d and Fig. 1e. This is different from active nematic flows confined in channels, which can exhibit net transport of the normally turbulent-like flows, but without a predefined flow polarity [17, 22]. In particular, we observe that active flows organize, on both sides of the frontal tip of each triangular obstacle, as a pair of counter rotating vortices, which pump the AN layer. This is clearly highlighted in Fig. 1f and g, where we represent the average vorticity field of the active flows.
Active nematic flows are know to be organized by self-propelling +1/2 defects, which form through spontaneous unbinding of opposite ±$\pm 1/2$ pairs [20], and are also injected into the system by the active boundary layers that form along the obstacle walls [32]. We have measured the time-averaged orientational field around triangular obstacles, both from experiments and numerical simulations (Fig. 1h), and we have computed the corresponding topological charge density field, which is correlated with the local presence of defects (Fig. 1i). We observe that the order parameter is maximal on the obstacle walls, where the nematic director adopts, on average, planar alignment, and a pattern of aligned ±$\pm 1/2$ defects emerges along the virtual flow channels, with +1/2 defects pushing the flow and $-1/2$ defects being advected by it (Fig. 1j).
Our observations are consistent with earlier simulations that reported a steady vortex lattice forming in AN confined in an array of astroid-shaped obstacles [33]. In our case, the flows emerging from counter-rotating vortex pairs leads to net transport in a direction that is predefined by the orientation of the obstacles in the array. Inside the virtual channels, flow speed is non monotonous, being higher between vortex pairs. Overall, this leads to a unidirectional pumping mechanism. (Fig. Active nematic pumps). A similar flow rectification effect has been recently proposed using numerical simulations with half-astroid-shaped obstacles [34], consistent with the expected ratchet effect in the presence of asymmetric obstacles.
@@
The reported pumping mechanism depends on the number, size, shape, and spatial distribution of the triangles in the array. A stable vortex lattice emerges only if at least a 2 by 2 square lattice of obstacles is present (Fig. Active nematic pumps). We consider this arrangement to be the minimal system to achieve active pumping. In Fig. 1, both in experiments and in numerical simulations, the size and spacing of the triangles are chosen to be equal to the average spacing between defects in the absence of obstacles (see early stages of Movie S1), $D_{def}$, which we have determined to be of the order of 70 $\mu$m in experiments and 56 lattice units in simulations.
@@
Using numerical simulations, we next study the effect of these two parameters. In Fig. 2a, we show the average normalized horizontal velocity as a function of the normalized size of the triangles, $H/D_{def}$, for a fixed spacing $L=2.3D_{def}$. Note that, in the absence of AN pumping effects, the flows are turbulent-like and therefore, the net velocity is zero in any direction. We observe a quick increase of the pumping efficiency with obstacle size, until $H\simeq D_{def}$, slowly decreasing for larger obstacles. We have performed simulations for different activity coefficients (related to the concentration of adenosine triphosphate in experiments), and find that the results remain invariant as the velocity and spatial dimensions are rescaled with the intrinsic velocity and active length scale in the absence of obstacles at each activity [35, 20].
@@
We next use simulations to explore, for a constant activity, the optimal geometry of the square lattice of triangles, in terms of their size and spacing. The resulting phase diagram, shown in Fig. 2b, indicates that the optimal arrangement corresponds to $H\simeq D_{def}$ and $L\simeq 2D_{def}$, which we have used in the experiments and simulations in Fig. 1, and will be used henceforth by default. Our experimental technique allows to fabricate more compact active pump arrangements, as shown in Fig. Active nematic pumps, thus demonstrating the capabilities for miniaturization of these devices.
@@ II.2 Active pump systems
After determining the optimal geometric features of basic AN pumps, we consider the design of more intricate microfluidic systems that include multiple AN pump units (Fig. Active nematic pumps). In particular, we have studied the performance of square lattices of triangles for different number of columns and rows (Fig. 3).
By keeping the number of rows fixed, we can study the effect of increasing the number of columns (Fig. 3a-d), which, in the language of hydraulics, is the equivalent of concatenating AN pumping units in series. We observe that the average flow speed increases with the number of columns (Fig. 3b), until reaching a maximum value at about 10 pumps in series (Fig. 3d). We also consider the concurrent profiles of the hydrostatic pressure within the active fluid along the flow channels, and we observe that it increases steadily with the number of pumping elements in series, even when the average speed has reached its saturation value (Fig. 3c and inset in Fig. 3d). In our active pump system, pressure gradients are positively correlated with the flow velocity. This is contrary to flow through a microfluidic channel or through a porous medium, where velocity is negatively correlated with the driving pressure gradients, even in the case of pressurized active fluids [36]. In classical hydraulics, a system of ideal pumps in series yields a flow velocity that does not change with the number of pumps and a pressure that builds up linearly with the number of elements [37]. Interestingly, our active pumps behave as ideal hydraulic pumps when the number of elements in series is larger than 10, as observed in Fig. 3d. The impossibility of a continuously increasing speed in a sequence of active nematic pumps can be justified by the fact that energy injection is finite. Indeed, as speed increases, viscous dissipation will eventually match energy injection and the input and output speed of subsequent pumping elements will remain the same.
@@
Figure 5: Microfluidic applications of active nematic pumping. a Experiments where the output of an active pump system is injected into a channel with smooth walls of width $L_{c}\simeq 3D_{def}$, with $D_{def}\simeq 70\mu$m. Triangular obstacles and channel boundaries are overlaid on the map of the velocity field. Scale bar 100 $\mu$m. b Simulation of an active pump system injecting into a channel of width $L_{c}=2.3D_{def}$ and activity $\zeta D_{def}/(v_{r}\eta)=7.84$. We perform simulations in a closed box of dimensions $\rm 54.9\,D_{def}\times 13.7\,D_{def}$, with no-slip boundary conditions on all walls. c Horizontal velocity (averaged over the central $50\,\mu$m along the channel) relative to $v_{r}$ for the experiments in a. d Velocity field along the central part of the channel, averaged in time and for the different initial conditions in simulations, for three different normalized channel widths $L_{c}/D_{def}$. e Experimental time-averaged velocity field in the presence of a closed circuit of elementary active pumps. Scale bar $100\mu$m (see also Movie Active nematic pumps). f Simulations of larger scale flows guided by a lattice of triangles. Counterclockwise flow circulation is guided by the active pump system. The system, with dimensions $\rm 36.6\,D_{def}\times 18.3\,D_{def}$ (larger than the region shown in the panel), is in a closed box with solid walls at the borders. See also movie Active nematic pumps. g and h are an example of pump reversal. Average relative horizontal velocity profile inside the flow channels corresponding to experiments (g) and simulations (h) for an array of triangular obstacles where new, larger obstacles facing in the opposite direction are imprinted on top of the original ones. In experiments, the square array has 70 $\mu$m spacing, triangles have an initial side length 27 $\mu$m and final side length 50 $\mu$m. Obstacles are created at t = 20 s and modified at t = 267 s. In simulations, array spacing is $2.3D_{def}$, triangles have an initial side length $0.8D_{def}$ and final side length $1.6D_{def}$. Simulations are performed on a 5 x 5 array with periodic boundary conditions.
Next, we analyze the performance of a lattice of active pumps when we increase the number of rows, keeping the number of columns fixed at 10, i.e., 9 pumps in series (Fig. 3e,f). In the language of hydraulics, this corresponds to changing the number of pumps in parallel, $N$. The output velocity, which does not change with $N$, is the one expected with 9 pumps in series. We will thus characterize each pump system with its average flow rate, which we find to be proportional to $N$, just as it would be for a system of ideal hydraulic pumps in parallel (Fig. 3f). We have also monitored the average pressure drop across the system, which we find to be constant, independent of $N$. These observations indicate that neighboring pumping elements along the same column operate independently, as a system of ideal pumps, despite the fact that the separation between adjacent flow channels is permeable. This favors not only net transport but also matter exchange, as we show in the next section.